OpenStudy (anonymous):

One number is two more than another number. If 4 times the larger number is subtracted from 5 times the smaller,the result is 10. Find both numbers.

OpenStudy (anonymous):

Numbers are x, 2+x. 5x - 4(2+x) = 10. Simplify to x-8=10. x=18, larger = 18+2=20. Answer is 18 and 20.